# Noncommutativepre-Poisson bialgebras and relative Rota-Baxter operators

### v1

@@ 1. Introduction
1.1. (Noncommutative) Poisson algebras and (noncommutative) pre-Poisson algebras
@@ 1.3. Outline of the paper
Let $A$ be a vector space with a bilinear product $\ast:A\otimes A\longrightarrow A$, $(A,\ast)$ is called a pre-Lie algebra if
@@
A dendriform algebra is a vector space $A$ together with two bilinear maps $\succ,\prec:A\otimes A\longrightarrow A$ such that the following equalities hold for all $x,y,z\in A$:
@@
Let $A$ be a 2-dimensional vector space with a basis $\{e_{1},e_{2}\}$. Define bilinear maps $\succ,\prec,\ast:A\otimes A\longrightarrow A$ respectively by (only non-zero multiplications are listed):
@@
Let $A$ be a 3-dimensional vector space with a basis $\{e_{1},e_{2},e_{3}\}$. Define binary products $\succ,\prec,\ast:A\otimes A\longrightarrow A$ respectively by (only non-zero multiplications are listed):
@@
Let $A$ and $V$ be vector spaces. For a linear map $f:A\longrightarrow\hbox{End}(V)$, define a linear map $f^{*}:A\longrightarrow\hbox{End}(V^{*})$ by $\langle f^{*}(x)u^{*},v\rangle=\langle u^{*},f(x)v\rangle$ for all $x\in A,u^{*}\in V^{*},v\in V$, where $\langle\ ,\ \rangle$ is the usual pairing between $V$ and $V^{*}$.
@@
    If $(V,l_{\succ},r_{\succ},l_{\prec},r_{\prec},l_{\ast},r_{\ast})$ is a quasi-representation of $(A,\succ,\prec,\ast)$, then both $(V,l_{\succ},r_{\prec},-l_{\ast})$ and $(V,l_{\cdot},r_{\cdot},r_{\ast}-l_{\ast},r_{\ast})$ are quasi-representations of the sub-adjacent noncommutative Poisson algebra $(A,\cdot,[\ ,\ ])$, where $r_{\cdot}=r_{\succ}+r_{\prec},~l_{\cdot}=l_{\succ}+l_{\prec}$.
@@
    $(A,L_{\succ},R_{\prec},-L_{\ast})$ is a representation of the sub-adjacent noncommutative Poisson algebra $(A,\cdot,[\ ,\ ])$ and $(A^{*},R_{\prec}^{*},L_{\succ}^{*},-L_{\ast}^{*})$ is a quasi-representations of $(A,\cdot,[\ ,\ ])$.
@@
2.2. Phase spaces of noncommutative Poisson algebras
@@
Let $(A,\succ,\prec,\ast)$ be the 2-dimensional noncommutative pre-Poisson algebra with a basis $\{e_{1},e_{2}\}$ given in Example 2.3. Define a nondegenerate skew-symmetric bilinear form $\omega$ on $A$ as follows: $\omega(e_{1},e_{2})=-\omega(e_{2},e_{1})=1$ and $\omega(e_{2},e_{2})=\omega(e_{1},e_{1})=0$. Then $(A,\succ,\prec,\ast,\omega)$ is a quadratic noncommutative pre-Poisson algebra.
@@
A Manin triple of noncommutative pre-Poisson algebras is a quadruple $(A,A_{1},A_{2},\omega)$, where $(A,\omega)$ is a quadratic noncommutative pre-Poisson algebra, $(A_{1},\succ_{1},\prec_{1},\ast_{1})$ and $(A_{2},\succ_{2},\prec_{2},\ast_{2})$ are noncommutative pre-Poisson subalgebras of $A$, such that
@@
Let $(A,A_{1},A_{2},\omega)$ be a Manin triple of noncommutative pre-Poisson algebras. Then the noncommutative pre-Poisson algebras $A$, $A_{1}$ and $A_{2}$ are coherent.
@@
which indicates that Eq. (6) holds. Thus, $A_{1}$ is coherent. Analogously, $A$ and $A_{2}$ are coherent. ∎
@@
    $(A\oplus A^{*},A,A^{*},\omega)$ is a Manin triple of noncommutative pre-Poisson algebras with the bilinear form given by Eq. (66) and the isotropic subalgebras are $A$ and $A^{*}$.
@@
(a) $\Longrightarrow$ (b) Assume that $(A\oplus A^{*},\cdot,[\ ,\ ],\omega)$ is a phase space of the noncommutative Poisson algebra $(A,\cdot_{A},[\ ,\ ]_{A})$. According to Proposition 2.15, there is a coherent noncommutative pre-Poisson algebra structure $(\succ,\prec,\ast)$ on $A\oplus A^{*}$ defined by Eqs. (64)-(65) such that $(A\oplus A^{*},\succ,\prec,\ast)$ is quadratic noncommutative pre-Poisson algebra. In view of Corollary 2.18, $(A,\succ|_{A},\prec|_{A},\ast|_{A})$ and $(A^{*},\succ|_{A^{*}},\prec|_{A^{*}},\ast|_{A^{*}})$ are subalgebras of $(A\oplus A^{*},\succ,\prec,\ast)$. It is obvious that $A$ and $A^{*}$ are isotropic. Thus, $(A\oplus A^{*},A,A^{*},\omega)$ is a Manin triple of noncommutative pre-Poisson algebras. Combining Proposition 4.2 [3], Theorem 4.1.5 [4] and Proposition 2.10, it follows directly that (b) $\Longleftrightarrow$ (c) $\Longleftrightarrow$ (d) $\Longrightarrow$ (a). ∎
@@ 3.1. Dendriform bialgebras and pre-Lie bialgebras
    A dendriform coalgebra is a triple $(A,\Delta_{\succ},\Delta_{\prec})$, where $A$ is a vector space and $\Delta_{\succ},\Delta_{\prec}:A\longrightarrow A\otimes A$ are linear maps such that the following conditions hold:
@@
    A pre-Lie coalgebra is a vector space $A$ together with a linear map $\delta:A\longrightarrow A\otimes A$ such that
@@ 3.2. Noncommutative pre-Poisson bialgebras
3.3. Coboundary noncommutative pre-Poisson bialgebras and the noncommutative pre-Poisson Yang-Baxter equation
@@
For a vector space $A$, the isomorphism $A\otimes A^{*}\simeq Hom(A^{*},A)$ identifies an element $r\in A\otimes A$ with a map $T_{r}:A^{*}\longrightarrow A$. Explicitly,
@@
Let $(V,l_{\succ},r_{\succ},l_{\prec},r_{\prec},l_{\ast},r_{\ast})$ be a representation of a coherent noncommutative pre-Poisson algebra satisfying Eqs. (18)-(20). Assume that $(V^{*},r_{\cdot}^{*},-l_{\prec}^{*},-r_{\succ}^{*},l_{\cdot}^{*},r_{\ast}^{*}-l_{\ast}^{*},r_{\ast}^{*})$ is the dual representation of $A$ given by Proposition 2.6. Denote $\hat{A}=A\ltimes V^{*}$ and suppose that $T:V\longrightarrow A$ is a linear map which is identifies an element in $\hat{A}\otimes\hat{A}$ through ($Hom(V,A)\simeq A\otimes V^{*}\subseteq\hat{A}\otimes\hat{A}$). Then $r=T+\tau(T)$ is a symmetric solution of the NPP-YBE in the coherent noncommutative pre-Poisson algebra $\hat{A}$ if and only if $T$ is an $\mathcal{O}$-operator on $(A,\succ,\prec,\ast)$ associated to $(V,l_{\succ},r_{\succ},l_{\prec},r_{\prec},l_{\ast},r_{\ast})$.
@@ 4.2. Factorizable noncommutative pre-Poisson bialgebras
Let $(A,\succ,\prec,\ast,\Delta_{\succ,r},\Delta_{\prec,r},\delta_{r})$ be a noncommutative pre-Poisson bialgebra. Assume that $\{e_{1},...,e_{n}\}$ is a basis of $A$ and $\{e^{*}_{1},...,e^{*}_{n}\}$ is the dual basis. Let
@@ 5.1. Relative Rota-Baxter operators and the NPP-YBE
Let $\omega$ be a non-degenerate bilinear form on a vector space $A$. Then there is an isomorphism $\omega^{\sharp}:A\longrightarrow A^{*}$ given by
@@
Let $(A,\succ,\prec,\ast)$ be a noncommutative pre-Poisson algebra and $\omega$ be a non-degenerate bilinear form on $A$. Then $(A,\succ,\prec,\ast,\omega)$ is a quadratic noncommutative pre-Poisson algebra if and only if the corresponding $r_{\omega}\in A\otimes A$ given by Eq. (142) is skew-symmetric and invariant.
@@
Let $A$ be a vector space and $\omega$ be a non-degenerate skew-symmetric bilinear form. Let $r\in A\otimes A$, $\lambda\in k$ and a linear map $P$ given by Eq. (143). Then $r$ satisfies

### v2

@@ 1. Introduction
$1.$ (Noncommutative) Poisson algebras and (noncommutative) pre-Poisson algebras
@@ 1.3. Outline of the paper
Let AA be a vector space with a bilinear product $\ast:A\otimes A\longrightarrow A$, $(A,\ast)$ is called a pre-Lie algebra if
@@
A dendriform algebra is a vector space AA together with two bilinear maps $\succ,\prec:A\otimes A\longrightarrow A$ such that the following equalities hold for all $x,y,z\in A$:
@@
Let AA be a 2-dimensional vector space with a basis $\{e_{1},e_{2}\}$. Define bilinear maps $\succ,\prec,\ast:A\otimes A\longrightarrow A$ respectively by (only non-zero multiplications are listed):
@@
Let AA be a 3-dimensional vector space with a basis $\{e_{1},e_{2},e_{3}\}$. Define binary products $\succ,\prec,\ast:A\otimes A\longrightarrow A$ respectively by (only non-zero multiplications are listed):
@@
Let AA and $V$ be vector spaces. For a linear map $f:A\longrightarrow\hbox{End}(V)$, define a linear map $f^{*}:A\longrightarrow\hbox{End}(V^{*})$ by $\langle f^{*}(x)u^{*},v\rangle=\langle u^{*},f(x)v\rangle$ for all $x\in A,u^{*}\in V^{*},v\in V$, where $\langle\ ,\ \rangle$ is the usual pairing between $V$ and $V^{*}$.
@@
    If $(V,l_{\succ},r_{\succ},l_{\prec},r_{\prec},l_{\ast},r_{\ast})$ is a quasi-representation of $(A,\succ,\prec,\ast)$, then both $(V,l_{\succ},r_{\prec},l_{\ast})$ and $(V,l_{\cdot},r_{\cdot},r_{\ast}-l_{\ast},r_{\ast})$ are quasi-representations of the sub-adjacent noncommutative Poisson algebra $(A,\cdot,[\ ,\ ])$, where $r_{\cdot}=r_{\succ}+r_{\prec},~l_{\cdot}=l_{\succ}+l_{\prec}$.
@@
    $(A,L_{\succ},R_{\prec},L_{\ast})$ is a representation of the sub-adjacent noncommutative Poisson algebra $(A,\cdot,[\ ,\ ])$ and $(A^{*},R_{\prec}^{*},L_{\succ}^{*},-L_{\ast}^{*})$ is a quasi-representations of $(A,\cdot,[\ ,\ ])$.
@@
$2.$ Phase spaces of noncommutative Poisson algebras
@@
Let $(A,\succ,\prec,\ast)$ be the 2-dimensional noncommutative pre-Poisson algebra with a basis $\{e_{1},e_{2}\}$ given in Example 2.3. Define a nondegenerate skew-symmetric bilinear form $\omega$ on AA as follows: $\omega(e_{1},e_{2})=-\omega(e_{2},e_{1})=1$ and $\omega(e_{2},e_{2})=\omega(e_{1},e_{1})=0$. Then $(A,\succ,\prec,\ast,\omega)$ is a quadratic noncommutative pre-Poisson algebra.
@@
A Manin triple of noncommutative pre-Poisson algebras is a quadruple $(A,A_{1},A_{2},\omega)$, where $(A,\omega)$ is a quadratic noncommutative pre-Poisson algebra, $(A_{1},\succ_{1},\prec_{1},\ast_{1})$ and $(A_{2},\succ_{2},\prec_{2},\ast_{2})$ are noncommutative pre-Poisson subalgebras of AA, such that
@@
Let $(A,A_{1},A_{2},\omega)$ be a Manin triple of noncommutative pre-Poisson algebras. Then the noncommutative pre-Poisson algebras AA, $A_{1}$ and $A_{2}$ are coherent.
@@
which indicates that Eq. (6) holds. Thus, $A_{1}$ is coherent. Analogously, AA and $A_{2}$ are coherent. ∎
@@
    $(A\oplus A^{*},A,A^{*},\omega)$ is a Manin triple of noncommutative pre-Poisson algebras with the bilinear form given by Eq. (66) and the isotropic subalgebras are AA and $A^{*}$.
@@
(a) $\Longrightarrow$ (b) Assume that $(A\oplus A^{*},\cdot,[\ ,\ ],\omega)$ is a phase space of the noncommutative Poisson algebra $(A,\cdot_{A},[\ ,\ ]_{A})$. According to Proposition 2.15, there is a coherent noncommutative pre-Poisson algebra structure $(\succ,\prec,\ast)$ on $A\oplus A^{*}$ defined by Eqs. (64)-(65) such that $(A\oplus A^{*},\succ,\prec,\ast)$ is quadratic noncommutative pre-Poisson algebra. In view of Corollary 2.18, $(A,\succ|_{A},\prec|_{A},\ast|_{A})$ and $(A^{*},\succ|_{A^{*}},\prec|_{A^{*}},\ast|_{A^{*}})$ are subalgebras of $(A\oplus A^{*},\succ,\prec,\ast)$. It is obvious that AA and $A^{*}$ are isotropic. Thus, $(A\oplus A^{*},A,A^{*},\omega)$ is a Manin triple of noncommutative pre-Poisson algebras. Combining Proposition 4.2 [3], Theorem 4.1.5 [4] and Proposition 2.10, it follows directly that (b) $\Longleftrightarrow$ (c) $\Longleftrightarrow$ (d) $\Longrightarrow$ (a). ∎
@@ 3.1. Dendriform bialgebras and pre-Lie bialgebras
    A dendriform coalgebra is a triple $(A,\Delta_{\succ},\Delta_{\prec})$, where AA is a vector space and $\Delta_{\succ},\Delta_{\prec}:A\longrightarrow A\otimes A$ are linear maps such that the following conditions hold:
@@
    A pre-Lie coalgebra is a vector space AA together with a linear map $\delta:A\longrightarrow A\otimes A$ such that
@@ 3.2. Noncommutative pre-Poisson bialgebras
$3.$ Coboundary noncommutative pre-Poisson bialgebras and the noncommutative pre-Poisson Yang-Baxter equation
@@
For a vector space AA, the isomorphism $A\otimes A^{*}\simeq Hom(A^{*},A)$ identifies an element $r\in A\otimes A$ with a map $T_{r}:A^{*}\longrightarrow A$. Explicitly,
@@
Let $(V,l_{\succ},r_{\succ},l_{\prec},r_{\prec},l_{\ast},r_{\ast})$ be a representation of a coherent noncommutative pre-Poisson algebra satisfying Eqs. (18)-(20). Assume that $(V^{*},r_{\cdot}^{*},-l_{\prec}^{*},-r_{\succ}^{*},l_{\cdot}^{*},r_{\ast}^{*}-l_{\ast}^{*},r_{\ast}^{*})$ is the dual representation of AA given by Proposition 2.6. Denote $\hat{A}=A\ltimes V^{*}$ and suppose that $T:V\longrightarrow A$ is a linear map which is identifies an element in $\hat{A}\otimes\hat{A}$ through ($Hom(V,A)\simeq A\otimes V^{*}\subseteq\hat{A}\otimes\hat{A}$). Then $r=T+\tau(T)$ is a symmetric solution of the NPP-YBE in the coherent noncommutative pre-Poisson algebra $\hat{A}$ if and only if $T$ is an $\mathcal{O}$-operator on $(A,\succ,\prec,\ast)$ associated to $(V,l_{\succ},r_{\succ},l_{\prec},r_{\prec},l_{\ast},r_{\ast})$.
@@ 4.2. Factorizable noncommutative pre-Poisson bialgebras
Let $(A,\succ,\prec,\ast,\Delta_{\succ,r},\Delta_{\prec,r},\delta_{r})$ be a noncommutative pre-Poisson bialgebra. Assume that $\{e_{1},...,e_{n}\}$ is a basis of AA and $\{e^{*}_{1},...,e^{*}_{n}\}$ is the dual basis. Let
@@ 5.1. Relative Rota-Baxter operators and the NPP-YBE
Let $\omega$ be a non-degenerate bilinear form on a vector space AA. Then there is an isomorphism $\omega^{\sharp}:A\longrightarrow A^{*}$ given by
@@
Let $(A,\succ,\prec,\ast)$ be a noncommutative pre-Poisson algebra and $\omega$ be a non-degenerate bilinear form on AA. Then $(A,\succ,\prec,\ast,\omega)$ is a quadratic noncommutative pre-Poisson algebra if and only if the corresponding $r_{\omega}\in A\otimes A$ given by Eq. (142) is skew-symmetric and invariant.
@@
Let AA be a vector space and $\omega$ be a non-degenerate skew-symmetric bilinear form. Let $r\in A\otimes A$, $\lambda\in k$ and a linear map $P$ given by Eq. (143). Then $r$ satisfies